# DB-Explore: Automated Database Exploration and Instruction Synthesis for Text-to-SQL

### v1

@@ Abstract
Recent text-to-SQL systems powered by large language models (LLMs) have demonstrated remarkable performance in translating natural language queries into SQL. However, these systems often struggle with complex database structures and domain-specific queries, as they primarily focus on enhancing logical reasoning and SQL syntax while overlooking the critical need for comprehensive database understanding. To address this limitation, we propose DB-Explore, a novel framework that systematically aligns LLMs with database knowledge through automated exploration and instruction synthesis. DB-Explore constructs database graphs to capture complex relational schemas, leverages GPT-4 to systematically mine structural patterns and semantic knowledge, and synthesizes instructions to distill this knowledge for efficient fine-tuning of LLMs. Our framework enables comprehensive database understanding through diverse sampling strategies and automated instruction generation, bridging the gap between database structures and language models. Experiments conducted on the SPIDER and BIRD benchmarks validate the effectiveness of DB-Explore, achieving an execution accuracy of 52.1% on BIRD and 84.0% on SPIDER. Notably, our open-source implementation, based on the Qwen2.5-coder-7B model, outperforms multiple GPT-4-driven text-to-SQL systems in comparative evaluations, and achieves near state-of-the-art performance with minimal computational cost.
@@ 2 Related Work
Early work in the field of Text-to-SQL generation primarily relied on rule-based methods to parse natural language queries and map them to valid SQL statements (Li and Jagadish, 2014; Popescu et al., 2003, 2004). More recently, LLMs such as GPT-4 (Achiam et al., 2023) have facilitated this task by leveraging their powerful text priors to generate SQL (Li et al., 2023b). Building upon this foundation, diverse technical approaches have emerged, such as RESD-SQL, DTS-SQL and RSL-SQL (Li et al., 2023a; Pourreza and Rafiei, 2024; Cao et al., 2024). These methods focus on enhancing schema selection and employ schema linking techniques to improve the identification of database schemas. Another line of research addresses complex query decomposition—DIN-SQL, PTD-SQL and QPL (Pourreza and Rafiei, 2023; Luo et al., 2024; Eyal et al., 2023) adopt stepwise problem decomposition to break down complex SQL generation into manageable subtasks. Candidate generation strategies emerge in approaches like MSC-SQL, SQL-PaLM, Distillery and XiYan (Gorti et al., 2024; Sun et al., 2023a; Maamari et al., 2024; Gao et al., 2024b), which boost accuracy through multi-candidate SQL generation and selection. DAIL-SQL (Gao et al., 2024a) introduces a skeleton-aware few-shot demonstration selection to enhance relevance of examples. MAC (Wang et al., 2023) and EPI-SQL (Liu and Tan, 2024) incorporate self-correction frameworks for SQL refinement.
Notably, recent open-source 7B/13B models (e.g., CodeLlama (Roziere et al., 2023), Qwen2 (Hui et al., 2024)) demonstrate competitive Text-to-SQL performance. These technical advances have catalyzed the emergence of multiple pre-training/fine-tuning frameworks leveraging large-scale synthetic data. CODES (Li et al., 2024a)) incrementally pre-trains the StarCoder (Li et al., 2023c) model using a corpus related to SQL and incorporates various external plugins to enhance Text-to-SQL performance. DTS-SQL (Pourreza and Rafiei, 2024) employs SFT on a dedicated schema linking model to reduce the input complexity of the model and improve performance. SENSE (Yang et al., 2024b) enhances the model’s preference on correct SQL by synthesizing strong data and weak data and direct preference optimization (DPO) on LLMs (Rafailov et al., 2024). ROUTE (Qin et al., 2024) unifies multi-task fine-tuning through task-specific data synthesis. However, these methodologies predominantly focus on enhancing LLMs’ direct SQL generation capabilities while overlooking comprehensive database understanding.
Our work aims to enhance the Text-to-SQL capabilities of LLMs by dynamic mining the knowledge underlying database structure and semantics. Through database exploration and data synthesis, we extract database knowledge and inject it into base model, reducing potential risks caused by insufficient knowledge of databases.
@@ 3.1 Database Graph Construction and Sampling
To enable more effective exploration and learning, we propose two complementary sampling strategies:
  1. 1.
    Random Walk Sampling: Starting from selected nodes, the walker iteratively traverses adjacent nodes via randomly chosen edges until reaching a predefined sampling size. This method ensures comprehensive schema coverage but may overlook high-priority relationships.
  2. 2.
    Empirical Weighted Sampling: Normalized edge weights guide traversal probabilities, biasing selection toward high-weight edges. This strategy emphasizes schema subgraphs aligned with user query patterns, enhancing instruction relevance.
Both strategies generate subgraphs as input for subsequent database multi-knowledge exploration, balancing semantic diversity via random walks and user intent alignment via weighted sampling.
@@ 3.2 Database Multi-Knowledge Exploration
We propose a three-pronged exploration and data synthesis framework to explore database by instruction synthesis while minimizing computational overhead: 1) Semantic Knowledge Extraction, 2) Structural Pattern Mining, and 3) Progressive Instruction Synthesis. The prompt templates shown in Figure 1 can be found in Appendix A.
@@ 3.2.1 Semantic Knowledge Extraction
Self-instruction has been shown to be an effective way to synthesize data (Wang et al., 2022). The instruction synthesis pipeline operates through a self-instruct framework initialized using manually-crafted template. We prompt the LLM with $S$ and randomly sampled seed instructions as demonstrations to synthesize novel queries $Q$ through schema-constrained synthesis. For instance, given columns "order_amount" and "customer_country", the model generates: "List customers from Canada with total orders exceeding 500" The generated ⟨$Q$, $S$⟩ pairs are subsequently feedback to the LLM to produce executable SQL responses $R$.
To achieve optimal semantic diversity while preserving logical validity, we implement temperature scaling ($\tau$=0.7) during generation. This technique yields paraphrased variants such as "Find Canadian clients spending over 500" while maintaining schema constraints. The iterative process progressively enriches the instruction pool through random sampling, thereby expanding coverage of potential user intents.
@@ 3.2.2 Structural Pattern Mining
As Figure 2 showed we first employ empirical weighted sampling to select schema subgraphs $S$, prioritizing high-frequency foreign-key edges "price.ID $=$ production.id". Then for each $S$, SQL statements $R$ are automatically generated by traversing all nodes using:
  • JOIN: Link tables via foreign keys
  • Filters: Derived from column data types
  • Aggregations: Applied to numerical columns
A representative SQL output for subgraph $S$:
"SELECT price.id FROM price INNER JOIN production ON price.id = production.id WHERE production.country = ’USA’ AND price.price < 15000"
Surviving SQL is then translated into natural language queries via back instruct (Shen et al., 2023), where the LLM rephrases R into natural language while preserving structural accuracy converting the above SQL to:
"List IDs of North American-manufactured vehicles under \$15,000"
Figure 2: Structural Pattern Mining Illustration.
@@ 3.2.3 Progressive Instruction Synthesis
Figure 3: Progressive Instruction Synthesis Illustration.
The evolution process initiates with a base query sampled from the DB Graph. Through iterative refinement: 1. Condition Augmentation: Schema-derived nodes (foreign keys, constraints) are recursively appended to the syntax tree via DB Graph-guided sampling. 2. Complexity Scaling: The LLM orchestrates tree expansion by injecting compound conditions such as nested WHERE clauses and multi-table JOIN while preserving syntactic validity. 3. Paraphrase Generation: Each evolved SQL tree is verbalized into diverse natural language expressions through template-guided LLM rewriting.
This progressive paradigm generates instruction pairs spanning atomic queries to multi-hop reasoning tasks, effectively simulating real-world query evolutionary pattern.
@@ 3.3 Instruction Filtering
  1. 1.
    Execution Verification:
    All generated SQL candidates R are executed against the target database. Queries are discarded based on three criteria:
    • Syntax errors
    • Invalid column/table references
    • Execution time exceeding 25 seconds
  2. 2.
    Consistency Verification:
[…]
    • Faithful translation between natural language intent and SQL logic
    • Adherence to schema constraints in S
[…]
@@ 3.4 Supervised Fine-Tuning and Inference
Schema Linking Model
Schema linking model trained to identify schema $S_{i}$ relevant to user queries $Q_{i}$. Given above $T={\{(Q_{i},D_{i})\}}$ , wherein the input prompt and target response generated is relevant schema $S_{i}$. The primary objective during supervised fine-tuning for schema linking model is to minimize the following loss function:
$Loss=-\frac{1}{|T|}\sum_{i=1}^{|T|}p(S_{i}|Q_{i},D_{i},M_{i})$ (1)
where $p$ represents the probability assigned to $S_{i}$ by the schema linking model. Integrating, $M_{i}$ represents the metadata of $D_{i}$ and $Q_{i}$.
SQL Generation Model
Conditioned on the schema subset $S_{i}$ from the linking model, this component maps $\textlangle Q_{i},S_{i}\textrangle$to executable SQL $R_{i}$. We then fine-tune the LLM while minimizing the following loss function:
$Loss=-\frac{1}{|T|}\sum_{i=1}^{|T|}p(R_{i}|Q_{i},S_{i},M_{i})$ (2)
Here $M_{i}$ represents the metadata of $S_{i}$ and $Q_{i}$.
@@ 4 Experiments
Ours: DB-Explore + Qwen2.5-Coder-7B 84.0 79.3 52.1 55.8
@@
(1) SPIDER The Spider dataset (Yu et al., 2018) is a large-scale, cross-domain benchmark for Text-to-SQL tasks, designed to evaluate the generalization ability of models across diverse database structures. It consists of 8,659 training samples, and 1,034 development samples, spanning 200 databases and 138 domains. Spider provides a robust challenge for models, emphasizing adaptability to unfamiliar database schemas. (2) BIRD The BIRD dataset (Li et al., 2023b) is a more challenging Text-to-SQL benchmark compared to SPIDER, featuring 95 real-world databases across 37 professional domains, with a total storage size of 33.4GB. BIRD emphasizes the challenges of handling large-scale, noisy, and incomplete database values, alongside the need for efficient SQL query generation
@@
In our experimental evaluation, we conduct comprehensive comparisons with state-of-the-art Text-to-SQL approaches across multiple technical paradigms. Our baselines can be categorized into two groups: 1) The prompting methods with closed-source LLMs MCS-SQL (Lee et al., 2025) and Chase-SQL (Guo et al., 2021), which enhance SQL generation accuracy through multi-prompt engineering and systematic comparison of candidate results. MAC-SQL (Wang et al., 2023) and MAG-SQL (Xie et al., 2024) employ multi-agent architectures for error correction and consensus building. TA-SQL (Qu et al., 2024) implement task alignment mechanisms to address LLMs’ hallucination. DIN-SQL (Pourreza and Rafiei, 2023) adopts task decomposition to handle complex queries through step-wise problem simplification, while DAIL-SQL (Gao et al., 2024a) selects high quality demonstration through skeleton-aware few-shot demonstration selection. 2) Fine-tuning with open-source LLMs 7B DTS-SQL (Pourreza and Rafiei, 2024) utilizes a multistage strategies based methods with two-stage fine-tuning and generation pipeline. SENSE (Yang et al., 2024b) introduces preference learning between strong and weak data in the Text-to-SQL domain. Codes (Li et al., 2024a) enhances its Text-to-SQL capabilities by incorporating a large amount of pretraining data and external plugins. On the other hand, ROUTE (Qin et al., 2024) introduces multi-task learning into the Text-to-SQL domain.
@@ 5.1 Main Results
Methods Spider
Dev-EX Dev-TS
Ours: DB-Explore + Qwen2.5-Coder-7B 84.0 79.3
- w/o Semantic Knowledge Extraction 82.9 78.4
- w/o Structural Pattern Mining 82.1 77.9
- w/o Progressive Instruction Synthesis 83.1 78.7
@@
Results on SPIDER
Table 1 presents the performance comparison between our method and baseline approaches on the SPIDER development set. The results demonstrate that prompt-based baselines achieve promising performance when augmented with GPT-4, while methods such as SENSE, CODES, and ROUTE bridge the performance gap between open-source models and GPT-4 through extensive data synthesis. In contrast, our database exploration mechanism effectively enhances open-source models performance with minimal computational cost. On the SPIDER benchmark, our approach achieves 84.0% EX, surpassing all 7B-parameter model-based methods except CODES, and outperforming certain GPT-4 based approaches such as DAIL-SQL and DIN-SQL.
Figure 4: Data volume analysis of different data synthesis methods on SPIDER, CODES only reports the storage size for pre-training data.
Results on BIRD
Our method achieves 52.1% EX with Qwen2.5-Coder-7B on BIRD development set, as shown in Table 1. However, in practice, our approach is not fully applicable to the BIRD dataset, as the discrepancy between training and evaluation database. Nevertheless, the data synthesis paradigm demonstrates foundational Text-to-SQL enhancement capabilities, achieving this performance with only 40MB augmented training samples. We obtained a result of 55.8% on the VES metric of the BIRD dataset. The VES metric effectively reflects the execution efficiency of SQL, highlighting the high efficiency of our data synthesis method. As our approach focuses on improving LLMs’ understanding of databases rather than handling complex logic, the synthesized data maintains a high degree of consistency with the database structure. This reduces redundant generation and unnecessary database operations, leading to significant progress on the VES metric.
Data Analysis
@@
Unlike existing approaches that focus on enhancing LLMs’ generalization capabilities for SQL generation, DB-Explore adopts a database-centric data synthesis paradigm that prioritizes mining database-specific knowledge. This fundamental distinction enables our method to achieve competitive performance with significantly reduced data requirements. As illustrated in Figure 4, our approach achieves the lowest computational cost among all data synthesis-based Text-to-SQL methods. Specifically, DB-Explore demonstrates superior data efficiency by attaining higher EX scores on the SPIDER benchmark than SENSE and ROUTE while using only 79% and 21% of their respective data volumes. Although our performance on SPIDER slightly trails CODES (84.0% vs. 85.4%), this comparison merits particular attention given that our supervised fine-tuning data constitutes merely 0.15% of the pre-training data required by CODES.

### v2

@@ Abstract
Recent text-to-SQL systems powered by large language models (LLMs) have demonstrated remarkable performance in translating natural language queries into SQL. However, these systems often struggle with complex database structures and domain-specific queries, as they primarily focus on enhancing logical reasoning and SQL syntax while overlooking the critical need for comprehensive database understanding. To address this limitation, we propose DB-Explore, a novel framework that systematically aligns LLMs with database knowledge through automated exploration and instruction synthesis. DB-Explore constructs database graphs to capture complex relational schemas, leverages GPT-4 to systematically mine structural patterns and semantic knowledge, and synthesizes instructions to distill this knowledge for efficient fine-tuning of LLMs. Our framework enables comprehensive database understanding through diverse sampling strategies and automated instruction generation, bridging the gap between database structures and language models. Experiments conducted on the SPIDER and BIRD benchmarks validate the effectiveness of DB-Explore, achieving an execution accuracy of 67.0% on BIRD and 87.8% on SPIDER. Notably, our open-source implementation based on Qwen2.5-Coder-7B achieves state-of-the-art results at minimal computational cost, outperforming several GPT-4-driven Text-to-SQL systems.
@@ 2 Related Work
Early Text-to-SQL systems relied on handcrafted, rule-based parsers to map natural language into SQL (Li and Jagadish, 2014; Popescu et al., 2003). The advent of large language models (LLMs) such as GPT-4 (Achiam et al., 2023) shifted focus to leveraging rich language priors for SQL generation (Li et al., 2023b). Building on this paradigm, schema-focused methods such as RESD-SQL, DTS-SQL, and RSL-SQL enhance schema linking and selection to ground queries in database structures (Li et al., 2023a; Pourreza and Rafiei, 2024; Cao et al., 2024), while decomposition-based techniques like DIN-SQL, PTD-SQL, and QPL break complex queries into simpler sub-tasks (Pourreza and Rafiei, 2023; Luo et al., 2024; Eyal et al., 2023). MSC-SQL, SQL-PaLM, Distillery, and XiYan improve robustness through multi-candidate generation and ranking (Gorti et al., 2024; Sun et al., 2023a; Maamari et al., 2024; Gao et al., 2024b), and skeleton-aware selection in DAIL-SQL refines few-shot demonstrations (Gao et al., 2024a). Self-correction frameworks in MAC and EPI-SQL further reduce errors via iterative feedback (Wang et al., 2023; Liu and Tan, 2024).
In parallel, competitive open-source LLMs have spurred fine-tuning frameworks that exploit synthetic or task-specific data. CODES incrementally pre-trains StarCoder on SQL corpora and integrates external plugins for enhanced performance (Li et al., 2024a), whereas DTS-SQL and SENSE employ schema-aware fine-tuning and direct preference optimization (DPO) (Rafailov et al., 2024) to improve accuracy (Pourreza and Rafiei, 2024; Yang et al., 2024b). ROUTE unifies multi-task tuning through targeted data synthesis (Qin et al., 2024). Recently, reinforcement learning (RL) methods have emerged as a promising direction in the Text-to-SQL field. SQL-o1 (Lyu et al., 2025) employs Monte Carlo Tree Search (MCTS) (Świechowski et al., 2023) to enhance the reasoning process during SQL generation. SQL-R1 (Ma et al., 2025) and Reasoning-SQL (Pourreza et al., 2025) leverage the GRPO (Liu et al., 2024) framework and incorporate domain-specific reward functions, improving the accuracy and robustness of Text-to-SQL systems. Despite these advances, most work remains focused on direct SQL generation, often overlooking holistic database understanding.
@@ 3.1 Database Graph Construction and Sampling
To enable more effective exploration and learning, we propose two complementary sampling strategies: 1) Random Walk Sampling: Starting from selected nodes, the walker iteratively traverses adjacent nodes via randomly chosen edges until reaching a predefined sampling size. This method ensures comprehensive schema coverage but may overlook high-priority relationships. 2) Empirical Weighted Sampling: Normalized edge weights guide traversal probabilities, biasing selection toward high-weight edges. This strategy emphasizes schema subgraphs aligned with user query patterns, enhancing instruction relevance. Both strategies generate subgraphs as input for subsequent database multi-knowledge exploration, balancing semantic diversity via random walks and user intent alignment via weighted sampling.
@@ 3.2 Database Multi-Knowledge Exploration
We propose a three-pronged exploration and data synthesis framework to explore database by instruction synthesis while minimizing computational overhead: 1) Semantic Knowledge Extraction, 2) Structural Pattern Mining, and 3) Progressive Instruction Synthesis.
@@ 3.2.1 Semantic Knowledge Extraction
Self-instruction has been shown to be an effective way to synthesize data (Wang et al., 2022). The instruction synthesis pipeline operates through a self-instruct framework initialized using manually-crafted template. We prompt the LLM with $S$ and randomly sampled seed instructions as demonstrations to synthesize novel queries $Q$ through schema-constrained synthesis. The generated ⟨$Q$, $S$⟩ pairs are subsequently feedback to the LLM to produce executable SQL responses $R$.
To achieve optimal semantic diversity while preserving logical validity, we implement temperature scaling ($\tau$=0.8) during generation. The iterative process progressively enriches the instruction pool through random sampling, thereby expanding coverage of potential user intents.
@@ 3.2.2 Structural Pattern Mining
As shown in Figure 1, we first apply empirical weighted sampling to select schema subgraphs $S$, giving priority to high-frequency foreign-key edges. For each selected subgraph, SQL statements $R$ are automatically generated by traversing all nodes in the subgraph, incorporating JOIN operations based on foreign keys, filter conditions derived from column data types, and aggregation functions applied to numerical columns. The resulting SQL statements are then translated into natural language queries using the back-instruct method (Shen et al., 2023), where an LLM rewrites each SQL query $R$ into a structurally faithful natural language instruction $Q$.
@@ 3.2.3 Progressive Instruction Synthesis
The synthesis process initiates with a base query sampled from the DB Graph. Through iterative refinement: 1. Condition Augmentation: Schema-derived nodes (foreign keys, constraints) are recursively appended to the syntax tree via DB Graph-guided sampling. 2. Complexity Scaling: The LLM orchestrates tree expansion by injecting compound conditions such as nested WHERE clauses and multi-table JOIN while preserving syntactic validity. 3. Instruction Rewriting: Each evolved SQL tree is verbalized into diverse natural language expressions through LLM rewriting.
Each iteration introduces new constraints and conditions to the instruction, progressively increasing its complexity. We set the number of iterations to 3, corresponding to user instructions of varying difficulty levels: simple, moderate, and challenging. This progressive paradigm generates instruction pairs ranging from atomic queries to multi-hop reasoning tasks, effectively simulating the evolutionary patterns of real-world user queries.
@@ 3.3 Instruction Filtering
All generated SQL candidates $R$ are executed against the target database, and queries are discarded if they contain syntax errors, reference invalid columns or tables, or exceed an execution time limit of 25 seconds. The surviving $\textlangle Q,R\textrangle$ pairs are then subjected to LLM-based consistency verification to ensure faithful alignment between the natural language intent and the SQL logic, as well as strict adherence to schema constraints defined in $S$. Validated instances are formatted as augmented tuples $\textlangle Q,D,R,S\textrangle$, where $D$ denotes the database context. These high-quality instances are reintegrated into the seed dataset, forming a self-improving feedback loop that drives iterative synthesis and database exploration.
@@ 4 Experiments
OmniSQL-7B (Li et al., 2025) 81.6 - 63.9 -
SQL-o1 + Qwen2.5-7B (Lyu et al., 2025) 84.7 78.5 66.7 70.4
SQL-R1(self-consistency@8) (Ma et al., 2025) 87.6 - 66.6 -
Reasoning-SQL + Qwen2.5-Coder-7B(Pourreza et al., 2025) - - 64.0 -
Ours: DB-Explore + Qwen2.5-Coder-7B 87.2 80.2 65.2 68.9
Ours: DB-Explore + Qwen2.5-Coder-7B + self-consistency@8 87.8 81.9 67.0 71.2
@@
(1) SPIDER The SPIDER dataset (Yu et al., 2018) is a large-scale, cross-domain benchmark for Text-to-SQL tasks, designed to evaluate the generalization ability of models across diverse database structures. It consists of 8,659 training samples, and 1,034 development samples, spanning 200 databases and 138 domains. (2) BIRD The BIRD dataset (Li et al., 2023b) is a more challenging Text-to-SQL benchmark compared to SPIDER, featuring 95 real-world databases across 37 professional domains, with a total storage size of 33.4GB.
@@
We evaluate DB-Explore against state-of-the-art Text-to-SQL methods across two major paradigms. The first includes approaches Prompting with Closed-Source LLMs, such as MCS-SQL (Lee et al., 2025) and Chase-SQL (Guo et al., 2021), which leverage multi-prompt strategies and result selection. MAC-SQL (Wang et al., 2023) and MAG-SQL (Xie et al., 2024) adopt multi-agent frameworks for error correction, while TA-SQL (Qu et al., 2024), DIN-SQL (Pourreza and Rafiei, 2023), and DAIL-SQL (Gao et al., 2024a) improve generation through task alignment, decomposition, and demonstration selection. The second group involves Fine-Tuning Open-Source 7B LLMs. DTS-SQL (Pourreza and Rafiei, 2024) employs a two-stage pipeline, SENSE (Yang et al., 2024b) leverages preference learning, and CODES (Li et al., 2024a) incorporates extensive pretraining and plugins. ROUTE (Qin et al., 2024) introduces multi-task learning, while SQL-o1 (Lyu et al., 2025), SQL-R1 (Ma et al., 2025), and Reasoning-SQL (Pourreza et al., 2025) integrate reinforcement learning to enhance SQL generation.
@@ 5.1 Main Results
Figure 2: Data volume analysis of different data synthesis methods on BIRD, CODES only reports the storage size for pre-training data
SPIDER BIRD
DB-Explore Base Model DB-Explore Base Model
Full-Schema EX 83.9 80.0 63.5 53.7
Filtered-Schema EX 87.2 - 65.2 -
Gold-Schema EX 89.1 83.4 69.9 63.8
Table 2: The performance of the SQL generation model for different schema inputs, where the filtered-schema is the result of the DB-Explore schema linking model.
Methods SPIDER BIRD
Dev-EX Dev-EX
Ours: DB-Explore + Qwen2.5-Coder-7B 87.2 65.2
- w/o Semantic Knowledge Extraction 86.1↓1.1 64.8↓0.4
- w/o Structural Pattern Mining 83.9↓3.3 62.5↓2.7
- w/o Progressive Instruction Synthesis 84.8↓2.4 62.2↓3.0
- w/o Instruction Filtering 83.7↓3.5 61.4↓3.8
@@
Performance on Main Benchmarks
As shown in Table 1, DB-Explore achieves strong performance, attaining execution accuracies of 87.8% on the SPIDER development set and 67.0% on the BIRD development set—setting a new state-of-the-art among methods based on open-source 7B models. Compared to the baseline using Qwen2.5-Coder-7B, our approach yields substantial improvements. Notably, DB-Explore outperforms reinforcement learning-based methods such as SQL-R1 (Ma et al., 2025) and Reasoning-SQL (Pourreza et al., 2025), despite relying only on limited supervised fine-tuning (SFT), significantly reducing training costs. Moreover, DB-Explore surpasses many Text-to-SQL methods that rely on closed-source models, including GPT-4 and GPT-4o. On the BIRD benchmark, it also achieves a VES score of 71.2%, underscoring that our database-specific exploration strategy can yield highly efficient query generation through capturing complex schema patterns.
Analysis of Data Volume
@@
Unlike existing approaches that focus on enhancing LLMs’ generalization capabilities for SQL generation, DB-Explore adopts a database-centric data synthesis paradigm that prioritizes mining database-specific knowledge. This fundamental distinction enables our method to achieve competitive performance with significantly reduced data requirements. As illustrated in Figure 2, our approach achieves the lowest computational cost among all data synthesis-based Text-to-SQL methods. Specifically, DB-Explore demonstrates superior data efficiency on BIRD benchmark, achieving a significantly higher EX accuracy than other data synthesis methods while using only 70.9% of the data used by ROUTE, and less than 1% of that used by OmniSQL and CODES.
Figure 3: DB-Explore performance trend under different database exploration ratios.
Analysis of DB Exploration Ratio
We analyze the impact of database exploration levels on model performance across the SPIDER and BIRD development sets. Different exploration levels are achieved by constraining the sampling scope of the DB Graph: 0% corresponds to the Qwen2.5-Coder-7B baseline without exploration, while 100% denotes full exploration as performed by DB-Explore. As shown in Figure 3, due to the relatively uniform distribution of queries across databases in both datasets, model performance increases almost linearly with the exploration ratio. Notably, database exploration not only enriches the model’s schema understanding but also enhances its overall Text-to-SQL capabilities through fine-tuning. As a result, performance gains are slightly more pronounced in the early stages of exploration compared to the later stages.
Figure 4: Performance comparison of DB-Explore at different instruction generation stages on BIRD development set.
Analysis of Progressive Instruction Synthesis
To investigate the effect of progressive instruction generation on DB-Explore’s ability to handle complex queries, we analyze its performance across different instruction synthesis stages on the BIRD development set. Each step corresponds to one iteration of evolving the training instructions; to prevent excessive distributional drift, we retain all simpler instructions from prior steps at every iteration. As shown in Figure 4, DB-Explore exhibits a significant improvement in handling simple queries at Step 1, where the synthesized instructions remain relatively basic. However, performance on moderate and challenging queries shows little change at this stage. In Steps 2 and 3, as the synthesized data becomes increasingly complex through iterative evolution, the model’s performance on moderate and challenging queries improves rapidly. As iterations accumulate, the synthesized instructions at Step 3, which now encompass queries involving four tables and five tables, are substantially more complex than those in the evaluation set. Although we retain all simpler instructions from Step 1 and Step 2, the overall data distribution at Step 3 still shifts slightly toward the more difficult examples, leading to a modest decline in performance on simple queries.
Analysis of Schema Linking
As shown in Table 2, the schema linking module not only alleviates the context length limitations of LLMs but also effectively focuses their attention on relevant schema elements. Incorporating Filtered-Schema leads to a 3.3% and 1.7% execution accuracy improvement on the SPIDER and BIRD datasets. Moreover, DB-Explore achieves strong performance when provided with the gold schema, further demonstrating its effectiveness on the Text-to-SQL task.